OpenStudy (anonymous):

68 – 18 ÷ 3 • 6 What's the answer?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

32 is correct

OpenStudy (anonymous):

68-18/3*6 18/3 = 6 *6 = 36 68-36=32

OpenStudy (anonymous):

ok I SEE

OpenStudy (anonymous):

explanation: you do multiplication and division first from left to right so 18/3=6 then 6*6=36 and then 68-36=32

OpenStudy (anonymous):

Yes so I was right lol

OpenStudy (anonymous):

yes you were xD

OpenStudy (anonymous):

thanks for double checking

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

have a good one